OpenStudy (anonymous):

The function f(x) = 800(1.09)4x shows the growth due to interest that Jim earns on an investment. The steps to calculate an equivalent function are shown below: [Step 1] f(x) = 800(1.09)4x [Step 2] f(x) = 800(1.094)x [Step 3] f(x) = 800(1.4116)x [Step 4] f(x) = (1129.27)x Which of the following is the first incorrect step?

OpenStudy (phi):

this is hard to read unless you put in ^ to show exponents.

OpenStudy (phi):

for example, I assume you mean [Step 1] f(x) = 800(1.09)^(4x) or using the equation editor \[ f(x) = 800(1.09)^{4x} \]?

OpenStudy (phi):

There is a rule for exponents : \[ \left(a^b\right)^c = a^{b\cdot c} \] that work either way. In other words, if you have the exponent 4*x then you can use the rule "backwards"